Accountancy

A business has earned average profit of rs 8000 during the last few years and the normal rate of return in similar type of business is 10% . Find out the value of goodwill by captalization method, give  that the assets of the firm amount to Rs 600000 and luabilities to Rs 100000
Answer #1

Capitalused value of average profits = 8000×100/10=800000

Goodwill=capitalused value - net assets 

= 800000-500000(600000-100000=300000

answered by: Advika